Find the equation of the cylinder with base π‘₯2 + 𝑦2 + 𝑧2 βˆ’ 3π‘₯ βˆ’ 6𝑧 + 9 = 0, π‘₯ βˆ’ 2𝑦 + 2𝑧 βˆ’ 6 = 0.


1
Expert's answer
2022-02-10T03:39:58-0500

Given the plane:

"\ud835\udc65 \u2212 2\ud835\udc66 + 2\ud835\udc67 \u2212 6 = 0"

and the sphere equation:

"\ud835\udc65^2 + \ud835\udc66^2 + \ud835\udc67^2 \u2212 3\ud835\udc65 \u2212 6\ud835\udc67 + 9 = 0"

"\displaystyle
\Rightarrow
\left(x-\frac{3}{2}\right)^2+\left(y-0\right)^2+(z-3)^2=\left(\frac{3}{2}\right)^2"by completing the square method

Then the center of the sphere is "\displaystyle
\left(\frac{3}{2},\ 0,\ 3\right)" with radius "\displaystyle
\frac{3}{2}."

Now, the length of the perpendicular from "\displaystyle
\left(\frac{3}{2},\ 0,\ 3\right)" to the plane "\ud835\udc65 \u2212 2\ud835\udc66 + 2\ud835\udc67 \u2212 6 = 0" is;

"\displaystyle
OM=\frac{|ax_1+by_1+cz_1+d|}{\sqrt{a^2+b^2+c^2}}"

Here, "\displaystyle
a=1, b=-2, c=2, d=-6, x_1=\frac{3}{2}, y_1=0, z_1=3". Thus,

"\displaystyle
OM=\frac{|1\cdot \frac{3}{2}+(-2)\cdot 0+2\cdot 3+(-6)|}{\sqrt{1^2+(-2)^2+2^2}}=\frac{|\frac{3}{2}|}{3}=\frac{1}{2}"

Next, "\displaystyle
AM^2=OA^2-OM^2=\left(\frac{3}{2}\right)^2-\left(\frac{1}{2}\right)^2=2"

The equation of the axis of the cylinder is;

"\displaystyle
\frac{x-\alpha}{l}=\frac{y-\beta}{m}=\frac{x-\gamma}{n}, \text{ Here, }(\alpha,\ \beta,\ \gamma)=\left(\frac{3}{2},\ 0,\ 3\right),\ l=1,\ m=-2,\ n=2".

"\displaystyle
\Rightarrow
\frac{x-\frac{3}{2}}{1}=\frac{y-0}{-2}=\frac{x-3}{2}, \text{ and }l^2+m^2+n^2=9"


The equation of the Right Circular Cylinder is;

"\displaystyle
[n(y-\beta)-m(z-\gamma)]^2+[l(z-\gamma)-n(x-\alpha)]^2+[m(x-\alpha)-l(y-\beta)]^2=AM^2(l^2+m^2+n^2)\\
\Rightarrow\\
\left[2(y-0)+2(z-3)\right]^2+\left[1(z-3)-2\left(x-\frac{3}{2}\right)\right]^2+\left[-2\left(x-\frac{3}{2}\right)-1(y-0)\right]^2=2(9)\\
\Rightarrow \left[2y+2z-6)\right]^2+\left[z-2x\right]^2+\left[-2x+3-y\right]^2=18\\
\Rightarrow 8x^2+5y^2+5z^2+8yz-4xz+4xy-12x-30y-24z+27=0"